# Topology and Spectrum in Measurement-Induced Phase Transitions

### v2

@@ Abstract
Competition among repetitive measurements of noncommuting observables and unitary dynamics can give rise to a rich variety of entanglement phases. We here characterize topological phases in monitored quantum systems by their spectrum and many-body topological invariants. We analyze (1+1)-dimensional monitored circuits for Majorana fermions, which have topological and trivial area-law entangled phases and a critical phase with sub-volume-law entanglement, through the Lyapunov spectrum. We uncover the presence (absence) of edge-localized zero modes inside the bulk gap in the topological (trivial) area-law phase and a bulk gapless spectrum in the critical phase. Furthermore, by suitably exploiting the fermion parity with twisted measurement outcomes at the boundary, we construct a topological invariant that sharply distinguishes the two area-law phases and dynamically characterizes the critical phase. Our work thus paves the way to extend the bulk-edge correspondence for topological phases from equilibrium to monitored quantum dynamics.
Introduction.–
@@
Topology is undoubtedly one of the most essential concepts in understanding stable phases of matter [1]. For ground states of local Hamiltonians with a finite excitation gap, states belonging to distinct topological phases cannot be smoothly deformed to each other without closing the gap, owing to discrete characters of their topological invariants [2]. In particular, for symmetry-protected topological phases [3, 4, 5, 6, 7, 8], nontrivial topology of the bulk leads to gapless boundary states robust against any symmetry-preserving perturbations. This phenomenon, known as the bulk-edge correspondence, has remarkable consequences in physical properties of the topological phases and has been intensively studied in the past few decades [9, 10, 11, 12, 13].
@@
Despite notable successes in equilibrium systems, topology in out-of-equilibrium systems caused by external environment has yet to be fully understood. In particular, it is still elusive how the topology plays a role in monitored quantum systems, which gather a recent extensive interest as they host novel non-equilibrium phases concerning, e.g., entanglement [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Indeed, repetitive measurements stabilizing distinct topological phases can lead to phase transitions between different entanglement phases, which can be probed by topological entanglement entropy or purification dynamics [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] (see also [43, 44, 45]). However, intrinsic spatiotemporal randomness caused by the probabilistic nature of measurement outcomes makes it difficult to generalize more standard notions of topological phases, such as bulk topological invariants and gapless edge modes, to monitored systems.
@@
Figure 1: (a) Monitored quantum circuit for Majorana fermions. In each time step, after unitary time evolution generated by ^$\hat{\mathcal{H}}_{\mathrm{Kitaev}}$ in Eq. (1), we weakly measure the Majoranas in the brickwork manner with Kraus operators (2). (b) Bulk topological invariant $\chi$ for the monitored system, obtained from the fermion parities $P^{\mathrm{PBC}/\mathrm{APBC}}$ for circuits under PBC/APBC. The APBC involves the boundary measurement operators with twisted outcomes for the PBC. (c) The measurement-induced phases characterized by the Lyapunov gap, existence of the edge mode, and our topological invariant. The topological (trivial) gapped phase exhibits the presence (absence) of edge mode and $\chi=-1$ ($\chi=+1$). The gapless critical phase is characterized by the dynamical behavior of $\chi$.
@@
In this Letter, we discover gapless edge modes and a bulk topological invariant that characterize measurement-induced phase transitions of monitored Majorana circuits and discuss their bulk-edge correspondence. Introducing effective Hamiltonians obtained through the Lyapunov analysis, we reveal the presence (absence) of edge-localized zero modes inside the bulk gap in the topological (trivial) area-law phase. We also construct a bulk topological invariant based on the fermion parity, extending the method in equilibrium systems to the monitored setting; this is accomplished by introducing a unique boundary condition with twisted measurement outcomes. The invariant sharply distinguishes these gapped area-law phases, featuring the bulk-edge correspondence. Moreover, we show that the critical entanglement phase corresponds to a gapless phase concerning the Lyapunov spectrum. While the bulk-edge correspondence is obscured in this phase, we demonstrate that the topological invariant can still dynamically characterize the two gapped and gapless phases. Our result is summarized in Fig. 1.
Model and Lyapunov analysis.–
@@
where $\gamma_{\ell}$ are Majorana fermions obeying $\{\gamma_{\ell},\gamma_{\ell^{\prime}}\}=2\delta_{\ell\ell^{\prime}}$ and $J,J^{\prime}$ are real constants. The measurements of Majorana pairs on odd/even bonds with an outcome $s=\pm 1$ are described by the Kraus operators,
@@
which are weak measurements of the strength $\theta_{e/o}=\tanh^{-1}(\mu_{e/o})$. We set $\mu_{e}=1-\mu_{o}$ hereafter. Given a (unnormalized) state |$\ket{\psi}$, each outcome $s$ is obtained with the Born probability, $p^{e/o}_{j}(s)=\bra{\psi}\hat{\mathcal{K}}_{j}^{e/o}(s)^{\dagger}\hat{\mathcal{K}}_{j}^{e/o}(s)\ket{\psi}/{\langle\psi|\psi\rangle}.$ In the following, we use three boundary conditions: open (OBC), periodic (PBC), and antiperiodic (APBC). For the unitary evolution $\hat{\mathcal{U}}$, the boundary condition is simply specified by $J^{\prime}$: $J^{\prime}=0,J,$ and $-J$ for the OBC, PBC, and APBC, respectively. For the measurements, the OBC means that we discard $\hat{\mathcal{K}}^{e}_{L}(s)$ from the circuit, while we need a careful definition of the APBC, as explained later.
At the time $t=T$, an initial state |$\ket{\psi_{0}}$ is evolved by the Kraus operator labeled by a sequence of outcomes $\bm{s}=\{\bm{s}_{1},\cdots,\bm{s}_{T}\}$ with $\bm{s}_{t}=\{s_{1,t},\cdots,s_{2L,t}\}$,
@@
Since both unitary evolution and measurements are bilinear in Majorana fermions, this circuit maps a fermionic Gaussian state to another Gaussian state [46, 47, 48, 49, 50, 51, 52]. A fermionic Gaussian state is completely characterized by the Majorana covariance matrix $\Gamma_{t}$ whose element reads $(\Gamma_{t})_{\ell\ell^{\prime}}=(i/2)\bra{\psi_{t}}[\gamma_{\ell},\gamma_{\ell^{\prime}}]\ket{\psi_{t}}/\langle\psi_{t}|\psi_{t}\rangle$. Time evolution of the (unnormalized) Gaussian state is encoded in the transformation of the Majorana fermions, $\vec{\gamma}\to\hat{\mathcal{K}}_{T}^{\dagger}(\bm{s})\vec{\gamma}(\hat{\mathcal{K}}_{T}^{\dagger}(\bm{s}))^{-1}=K_{T}(\bm{s})\vec{\gamma}$, where
@@
Here, $H_{\mathrm{Kitaev}}$ and $\Theta^{e/o}(\bm{s}_{t})$ are real antisymmetric matrices defined through ^$\hat{\mathcal{H}}_{\mathrm{Kitaev}}=i\vec{\gamma}^{\mathsf{T}}H_{\mathrm{Kitaev}}\vec{\gamma}/4$ and $\prod_{j}\hat{\mathcal{K}}^{e/o}_{j}(s_{2j/2j-1,t})\propto\exp({i\vec{\gamma}^{\mathsf{T}}\Theta^{e/o}(\bm{s}_{t})}\vec{\gamma}/4)$ with $\vec{\gamma}=(\gamma_{1},\cdots,\gamma_{2L})^{\mathsf{T}}$ (see Supplemental Material [53]).
Since this circuit $K_{T}(\bm{s})$ is a product of random matrices, we can perform the Lyapunov analysis [54, 55, 56, 57, 58, 59, 60]. Provided that the Oseledec theorem [61, 62, 63] holds, the Lyapunov spectrum $z_{\ell}$ does not depend on the sequence of outcomes $\bm{s}$. It is also related to the energy spectrum of the effective Hamiltonian,
@@
and the corresponding many-body Hamiltonian ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})=-i\vec{\gamma}^{\mathsf{T}}H_{\textrm{eff},T}(\bm{s})\vec{\gamma}/4$. Since $H_{\textrm{eff},T}(\bm{s})$ is real antisymmetric, its eigenvalues $iz_{\ell,T}(\bm{s})\in i\mathbb{R}$ come in pairs, $z_{2j-1,T}(\bm{s})=-z_{2j,T}(\bm{s})$. This defines the single-particle energy spectrum $z_{\ell,T}(\bm{s})$, which in the asymptotic limit coincides with the Lyapunov spectrum: $z_{\ell}=\lim_{T\to\infty}z_{\ell,T}(\bm{s})$. We below arrange them as $0\leq z_{1}=-z_{2}\leq\cdots\leq z_{2L-1}=-z_{2L}$. We can also compute the corresponding Lyapunov vectors $\vec{w}_{\ell,T}(\bm{s})$. Written in the matrix form $\widetilde{O}_{T}(\bm{s})=(\vec{w}_{1,T}(\bm{s}),\cdots,\vec{w}_{2L,T}(\bm{s}))$, they give in the asymptotic limit an orthogonal matrix that brings $H_{\textrm{eff},T}(\bm{s})$ into the standard form,
@@
In practice, we construct $\widetilde{O}_{T}(\bm{s})$ through computing Lyapunov vectors corresponding to the non-negative Lyapunov spectrum based on the complex fermion representation, which is equivalent to the Majorana representation [53]. In the following, the initial state of the trajectories is fixed to be a vacuum state |$\ket{\psi_{0}}=\ket{0}$, whereas the Lyapunov analysis itself is performed with random initial vectors $\vec{w}_{\ell,0}$ for a given trajectory. We only look at a single trajectory specified by a typical sequence of $\bm{s}$ and consider a temporal average of quantities in the long-time limit unless otherwise mentioned [53].
@@
Figure 2: (a) Non-negative single-particle Lyapunov spectra for $L=64$ and (b) sum of squared amplitudes of the spinor wavefunction corresponding to the lowest energy $z_{1}$ at the edges for the measurement-only circuit with the OBC. We find that the topologically non-trivial phase hosts a spatially localized Majorana edge mode with the zero Lyapunov exponent. The inset in (b) shows the spatial distribution of squared amplitudes of the spinor wavefunction at $t=2L$ averaged over $1000$ different trajectories for $L=16$ and $\mu_{e}=0.1$. (c,d) Lowest non-negative single-particle Lyapunov spectrum for the measurement-only circuit with (c) PBC and (d) APBC. They are gapped except for $\mu_{e}\simeq 0.5$, where $z_{1}$ behaves as $\sim 1/L$ for PBC and the exact gap closing is anticipated for APBC. (e) Topological invariant after a sufficiently long time, which sharply characterizes topological and trivial phases.
Lyapunov edge modes in measurement-only circuits.–
@@
We first perform the Lyapunov analysis for the measurement-only circuit ($J=J^{\prime}=0$). In this case, a direct transition at $\mu_{e}\simeq 0.5$ occurs from a topological to a trivial area-law phase, as numerically checked by the topological entanglement entropy and bipartite mutual information [53], as done in [41, 34, 42].
@@
Remarkably, we reveal that the topological area-law phase is characterized by Majorana zero modes for the Lyapunov spectrum. Figure 2(a) shows the non-negative single-particle Lyapunov spectrum $z_{k}$ against $\mu_{e}$ for the measurement-only circuit under the OBC for $L=64$. Importantly, $z_{1}$ takes a finite value for $\mu_{e}>0.5$, while it vanishes for $\mu_{e}\leq 0.5$. Vanishing of $z_{1}$ under the OBC in the topological area-law phase implies the many-body Lyapunov gap closing for ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})$, which is reminiscent of Majorana zero modes for the static Kitaev chain [64].
@@
We next show that the above zero mode is a spatially localized edge mode by analyzing the spinor wavefunction $\vec{\psi}_{T}(\bm{s})$ [65] with its element $(\vec{\psi}_{T}(\bm{s}))_{j}=\sqrt{\sum_{a=1,2}\sum_{b=2j-1,2j}((\vec{w}_{a,T}(\bm{s}))_{b}^{2}/2}$ corresponding to the lowest energy $z_{1,T}(\bm{s})\simeq z_{1}$. In Fig. 2(b), We plot the long-time average of the sum of the squared amplitudes for $\vec{\psi}_{T}(\bm{s})$ at both edges, $|(\vec{\psi}_{T}(\bm{s}))_{1}|^{2}+|(\vec{\psi}_{T}(\bm{s}))_{L}|^{2}$, against $\mu_{e}$. Deep inside the topological phase ($\mu_{e}\ll 0.5$), it takes a value close to $1$ regardless of the system size, signaling localization of the zero mode near the edges. In contrast, it decreases toward a value $2/L$ in the trivial phase, indicating that the Majorana mode delocalizes over the whole lattice on average. The localization of the zero modes in the topological phase is much more visible if we directly look at the spatial profile of the $|(\vec{\psi}_{T}(\bm{s}))_{j}|^{2}$. The inset of Fig. 2(b) shows (trajectory-averaged) values of $|(\vec{\psi}_{T}(\bm{s}))_{j}|^{2}$ for $L=16$ and $\mu_{e}=0.1$. These results indicate that the Lyapunov spectrum clearly reveals edge-localized Majorana zero modes in the topological area-law phase.
Topological invariant and bluk-edge correspondence.–
@@
${Q_{T}(\bm{s})=\mathrm{sgn}(\mathrm{Pf}[{H}_{\mathrm{eff},T}^{\mathrm{PBC}}(\bm{s})]\mathrm{Pf}[{H}_{\mathrm{eff},T}^{\mathrm{APBC}}(\bm{s})])}.$ (7)
However, the explicit form of $H_{\textrm{eff},T}(\bm{s})$ is numerically intractable for large $T$. Thus, we instead compute $\chi_{T}(\bm{s})$ defined by
@@
Since $\chi_{T}(\bm{s})$ approximates and converges to $Q_{T}(\bm{s})$ for $T\to\infty$, we will also call $\chi_{T}(\bm{s})$ as the topological invariant. We note that $\widetilde{O}^{\textrm{PBC/APBC}}_{T}(\bm{s})$ is not an orthogonal matrix for general $T$.
In Fig. 2(c,d), we show the lowest single-particle Lyapunov spectrum $z_{1}$ for the measurement-only circuit under the (c) PBC and (d) APBC. It takes a finite value for both cases, implying a finite bulk Lyapunov gap, within each area-law phase. As shown in Fig. 2(e), the topological invariant $\chi_{T}(\bm{s})$ clearly separates the topological and trivial area-law phases. Note that these results are qualitatively independent of the measurement outcomes $\bm{s}$. In fact, $P^{\mathrm{PBC}}_{T}(\bm{s})=1$ is satisfied in the whole parameter region, since the gap does not close at the transition $\mu_{e}\simeq 0.5$ for PBC due to finite-size splitting $z_{1}\sim 1/L$. On the other hand, $P^{\mathrm{APBC}}_{T}(\bm{s})$ abruptly changes from $-1$ to $+1$ across the transition point for APBC, implying an exact gap closing near $\mu_{e}=0.5$. This results in the observed transition in $\chi_{T}(\bm{s})$. The behaviors of the bulk Lyapunov gap resemble those of the bulk energy gap $\Delta E$ of the static translation-invariant Kitaev chain [64]; in that case, $\Delta E=0$ for PBC and $\Delta E\sim 1/L$ for APBC at the transition, across which the Pfaffian invariant changes its sign.
The edge-localized zero mode under the OBC in Fig. 2(a,b) and the topological invariant in Fig. 2(e) indicate that the bulk-edge correspondence applies even to nonequilibrium phases under temporally random measurements, if the bulk gap is open. Such an exploration based on topology in random dynamics is accomplished by the Lyapunov analysis and the appropriate definition of the topological invariant via effective Hamiltonians. Note that a naive quantity $P_{T}^{\textrm{PBC}}(\bm{s})P^{\textrm{APBC}}_{T}(\bm{s}^{\prime})$ with $\bm{s}\neq\bm{s}^{\prime}$ does not provide a good topological invariant in general. Rather, focusing on the trajectory under the APBC dependent on the trajectory under PBC makes their fermion parities meaningful to define the invariant under random measurements.
Gapless critical phase with additional unitary.–
Figure 3: (a) Non-negative single-particle Lyapunov spectrum for the monitored circuit with $J=0.5$ under the OBC for $L=64$. (b,c) Lowest non-negative spectrum $z_{1}$ under the (b) PBC and (c) APBC. (d) Topological invariant at $t=L$ averaged over $1000$ different trajectories. The Lyapunov spectrum and topological invariant distinguish the topological area-law, critical gapless, and trivial area-law phases at $\mu_{e}\simeq 0.4$ and $0.6$, as indicated by the dashed vertical lines.

### v4

@@ Abstract
Competition among repetitive measurements of noncommuting observables and unitary dynamics can give rise to a wide variety of entanglement phases. Here, we propose a general framework based on Lyapunov analysis to characterize topological properties in monitored quantum systems through their spectrum and many-body topological invariants. We illustrate this framework by analyzing (1+1)-dimensional monitored circuits for Majorana fermions, which are known to exhibit topological and trivial area-law entangled phases as well as a critical phase with sub-volume-law entanglement. Through the Lyapunov analysis, we identify the presence (absence) of edge-localized zero modes inside the bulk gap in the topological (trivial) phase and a bulk gapless spectrum in the critical phase. Furthermore, by suitably exploiting the fermion parity with twisted measurement outcomes at the boundary, we construct a topological invariant that distinguishes the two area-law phases and dynamically characterizes the critical phase. Our framework thus provides a general route to extend the notion of bulk–edge correspondence to monitored quantum dynamics.
Introduction–
@@
Topology is undoubtedly one of the most essential concepts in understanding stable phases of matter [1]. For the ground states of local Hamiltonians with a finite excitation gap, states belonging to distinct topological phases cannot be smoothly deformed to each other without closing the gap, owing to discrete characters of their topological invariants [2]. In particular, for symmetry-protected topological phases [3, 4, 5, 6, 7, 8], the nontrivial topology of the bulk leads to gapless boundary states robust against any symmetry-preserving perturbations. This phenomenon, known as the bulk-edge correspondence, has remarkable consequences in the physical properties of the topological phases and has been intensively studied in the past few decades [9, 10, 11, 12, 13].
@@
Despite notable successes in equilibrium systems, topology in out-of-equilibrium systems caused by the external environment has yet to be fully understood. In particular, it is still elusive how the topology plays a role in monitored quantum systems, which have garnered significant interest recently as they host novel nonequilibrium phases concerning, e.g., entanglement [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Indeed, repetitive measurements stabilizing distinct topological phases can lead to phase transitions between different entanglement phases, which can be probed by topological entanglement entropy or purification dynamics [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] (see also [43, 44, 45]). However, intrinsic spatiotemporal randomness caused by the probabilistic nature of measurement outcomes makes it difficult to generalize more standard notions of topological phases, such as bulk topological invariants and gapless edge modes, to monitored systems.
@@
Figure 1: (a) Monitored quantum circuit for Majorana fermions. In each time step, after unitary time evolution generated by ^$\hat{\mathcal{H}}_{\mathrm{Kitaev}}$ in Eq. (1), we weakly measure the Majoranas in the brickwork manner with Kraus operators in Eq. (2). (b) Bulk topological invariant $\chi$ for the monitored system, obtained from the fermion parities $P^{\mathrm{PBC}/\mathrm{APBC}}$ for circuits under PBC/APBC. The APBC involves the boundary measurement operators with twisted outcomes with respect to the PBC. (c) The measurement-induced phases characterized by the Lyapunov gap, the existence of the edge modes, and our topological invariant. The topological (trivial) gapped phase exhibits the presence (absence) of edge modes and $\chi=-1$ ($\chi=+1$). The critical gapless phase is characterized by the dynamical behavior of $\chi$.
@@
In this Letter, we discover gapless edge modes and a bulk topological invariant that characterize measurement-induced phase transitions of monitored Majorana circuits and discuss their bulk-edge correspondence. Introducing effective Hamiltonians obtained through the Lyapunov analysis, we reveal the presence (absence) of edge-localized zero modes inside the bulk gap in the topological (trivial) area-law phase. We also construct a bulk topological invariant based on the fermion parity, extending the method in equilibrium systems to the monitored setting; this is accomplished by introducing a unique boundary condition with twisted measurement outcomes. The invariant sharply distinguishes these gapped area-law phases, featuring the bulk-edge correspondence. Moreover, we show that the critical entanglement phase corresponds to a gapless phase concerning the Lyapunov spectrum. While the bulk-edge correspondence is obscured in this phase, we demonstrate that the topological invariant can still dynamically characterize the two gapped and gapless phases. While we mainly focus on free fermionic systems in this Letter, our methodologies can be extended to many-body interesting systems and thus offer a general framework for the study of measurement-induced topological phases. Our result is summarized in Fig. 1.
Model and Lyapunov analysis–
@@
where $\gamma_{\ell}$ are Majorana fermions obeying $\{\gamma_{\ell},\gamma_{\ell^{\prime}}\}=2\delta_{\ell\ell^{\prime}}$ and $J,J^{\prime}$ are real constants. The measurements of parities for neighboring Majorana pairs $i\gamma_{\ell}\gamma_{\ell+1}$ on odd or even bonds with an outcome $s=\pm 1$ are described by the Kraus operators [46] ,
@@
which are weak measurements of the strength $\theta_{e/o}=\tanh^{-1}(\mu_{e/o})$. The Kraus operator reduces to a projective measurement for $\theta_{o/e}\rightarrow\infty$ ($\mu_{o/e}\rightarrow 1$) or to the identity operation for $\theta_{o/e}\rightarrow 0$ ($\mu_{o/e}\rightarrow 0$). Experimental setups for implementing the Majorana parity measurements are discussed in [41, 47]. We set $\mu_{e}=1-\mu_{o}$ hereafter to simply interpolate between the limits $(\mu_{o},\mu_{e})=(0,1)$ and $(1,0)$, where the two distinct area-law entangled phases are certainly realized. Given a (un-normalized) state |$\ket{\psi}$, each outcome $s$ is obtained with the Born probability, $p^{e/o}_{j}(s)=\bra{\psi}\hat{\mathcal{K}}_{j}^{e/o}(s)^{\dagger}\hat{\mathcal{K}}_{j}^{e/o}(s)\ket{\psi}/{\langle\psi|\psi\rangle}.$ In the following, we use three boundary conditions: open (OBC), periodic (PBC), and antiperiodic (APBC). For the unitary evolution $\hat{\mathcal{U}}$, the boundary condition is simply specified by $J^{\prime}$: $J^{\prime}=0,J,$ and $-J$ for the OBC, PBC, and APBC, respectively. For the measurements, the OBC means that we discard $\hat{\mathcal{K}}^{e}_{L}(s)$ from the circuit, while we need a careful definition of the APBC, as explained later.
At the time $t=T$, an initial state |$\ket{\psi_{0}}$ is evolved by the Kraus operator labeled by a sequence of outcomes $\bm{s}=\{\bm{s}_{1},\ldots,\bm{s}_{T}\}$ with $\bm{s}_{t}=\{s_{1,t},\ldots,s_{2L,t}\}$,
@@
Since both unitary evolution and measurements are bilinear in Majorana fermions, this circuit maps a fermionic Gaussian state to another Gaussian state [48, 49, 50, 51, 52, 53, 54]. A fermionic Gaussian state is completely characterized by the Majorana covariance matrix $\Gamma_{t}$ whose element reads $(\Gamma_{t})_{\ell\ell^{\prime}}=(i/2)\bra{\psi_{t}}[\gamma_{\ell},\gamma_{\ell^{\prime}}]\ket{\psi_{t}}/\langle\psi_{t}|\psi_{t}\rangle$. Time evolution of the (un-normalized) Gaussian state is encoded in the transformation of the Majorana fermions, $\vec{\gamma}\to\hat{\mathcal{K}}_{T}^{\dagger}(\bm{s})\vec{\gamma}[\hat{\mathcal{K}}_{T}^{\dagger}(\bm{s})]^{-1}=K_{T}(\bm{s})\vec{\gamma}$, where
@@
Here, $H_{\mathrm{Kitaev}}$ and $\Theta^{e/o}(\bm{s}_{t})$ are real antisymmetric matrices defined through ^$\hat{\mathcal{H}}_{\mathrm{Kitaev}}=i\vec{\gamma}^{\mathsf{T}}H_{\mathrm{Kitaev}}\vec{\gamma}/4$ and $\prod_{j}\hat{\mathcal{K}}^{e/o}_{j}(s_{2j/2j-1,t})\propto\exp[{i\vec{\gamma}^{\mathsf{T}}\Theta^{e/o}(\bm{s}_{t})}\vec{\gamma}/4]$ with $\vec{\gamma}=(\gamma_{1},\ldots,\gamma_{2L})^{\mathsf{T}}$ (see Supplemental Material [55]).
Since this circuit $K_{T}(\bm{s})$ is a product of random matrices, we can perform the Lyapunov analysis [56, 57, 58, 59, 60, 61, 62]. Provided that Oseledec’s theorem [63, 64, 65] holds, the Lyapunov spectrum $z_{\ell}$ does not depend on the sequence of outcomes $\bm{s}$. It is also related to the energy spectrum of the effective Hamiltonian,
@@
and the corresponding many-body Hamiltonian ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})=-i\vec{\gamma}^{\mathsf{T}}H_{\textrm{eff},T}(\bm{s})\vec{\gamma}/4$. Since $H_{\textrm{eff},T}(\bm{s})$ is real antisymmetric, its eigenvalues $iz_{\ell,T}(\bm{s})\in i\mathbb{R}$ come in pairs, $z_{2j-1,T}(\bm{s})=-z_{2j,T}(\bm{s})$. This defines the single-particle energy spectrum $z_{\ell,T}(\bm{s})$, which in the asymptotic limit coincides with the Lyapunov spectrum: $z_{\ell}=\lim_{T\to\infty}z_{\ell,T}(\bm{s})$. We below arrange them as $0\leq z_{1}=-z_{2}\leq\cdots\leq z_{2L-1}=-z_{2L}$. We can also compute the corresponding Lyapunov vectors $\vec{w}_{\ell,T}(\bm{s})$. Written in the matrix form $\widetilde{O}_{T}(\bm{s})=(\vec{w}_{1,T}(\bm{s}),\ldots,\vec{w}_{2L,T}(\bm{s}))$, they give in the asymptotic limit an orthogonal matrix that brings $H_{\textrm{eff},T}(\bm{s})$ into the standard form,
@@
In practice, we construct $\widetilde{O}_{T}(\bm{s})$ through computing Lyapunov vectors corresponding to the non-negative Lyapunov spectrum based on the complex fermion representation, which is equivalent to the Majorana representation [55]. In the following, the initial state of the trajectories is fixed to be a vacuum state |$\ket{\psi_{0}}=\ket{0}$, whereas the Lyapunov analysis itself is performed with random initial vectors $\vec{w}_{\ell,0}$ for a given trajectory. We only look at a single trajectory specified by a typical sequence of $\bm{s}$ and consider a temporal average of quantities in the long-time limit unless otherwise mentioned [55]. Note that, in contrast to the conventional setup of products of random matrices, the Born rule unique to monitored systems makes the existence or convergence of the Lyapunov spectrum a mathematically nontrivial problem, which has only recently been addressed in Refs. [66, 67].
@@
Figure 2: (a) Non-negative single-particle Lyapunov spectra for $L=128$ and (b) sum of the squared amplitudes of the spinor wave function corresponding to the lowest energy $z_{1}$ at the edges for the measurement-only circuit with OBC. The inset in (b) shows the spatial distribution of the squared amplitudes of the spinor wave function at $t=2L$ averaged over $1000$ different trajectories for $L=16$ and $\mu_{o}=0.1$. (c),(d) Lowest non-negative single-particle Lyapunov spectrum for the measurement-only circuit with (c) PBC and (d) APBC. (e) Topological invariant after a sufficiently long time.
Lyapunov edge modes in measurement-only circuits–
@@
We first perform the Lyapunov analysis for the measurement-only circuit ($J=J^{\prime}=0$). In this case, a direct transition at $\mu_{o}\simeq 0.5$ occurs from a topological to a trivial area-law phase, as numerically checked by the topological entanglement entropy and bipartite mutual information [55], following the approach in [41, 34, 42].
@@
We find that the topological transition between the two area-law phases can be viewed as the transition between two bulk-gapped phases for the Lyapunov spectrum, and the bulk gap closes once at the transition. Remarkably, we reveal that the topological (trivial) area-law phase is distinguished by the presence (absence) of Majorana zero modes within the bulk gap. Figure 2(a) shows the non-negative single-particle Lyapunov spectrum $z_{k}$ against $\mu_{o}$ for the measurement-only circuit under the OBC for $L=128$. Importantly, $z_{1}$ takes a finite value for $\mu_{o}>0.5$, while it vanishes for $\mu_{o}\leq 0.5$. Vanishing of $z_{1}$ under the OBC in the topological area-law phase implies the many-body Lyapunov gap closing for ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})$, which is reminiscent of the appearance of Majorana zero modes for the static Kitaev chain [68].
@@
We next show that the above zero modes are spatially localized edge modes by analyzing the spinor wave function $\vec{\psi}_{T}(\bm{s})$ [69] with its element $[\vec{\psi}_{T}(\bm{s})]_{j}=\sqrt{\sum_{a=1,2}\sum_{b=2j-1,2j}[\vec{w}_{a,T}(\bm{s})]_{b}^{2}/2}$ corresponding to the lowest energy $z_{1,T}(\bm{s})\simeq z_{1}$. In Fig. 2(b), we plot the long-time average of the sum of the squared amplitudes for $\vec{\psi}_{T}(\bm{s})$ at both edges, $|[\vec{\psi}_{T}(\bm{s})]_{1}|^{2}+|[\vec{\psi}_{T}(\bm{s})]_{L}|^{2}$, against $\mu_{o}$. Deep inside the topological phase ($\mu_{o}\ll 0.5$), it takes a value close to $1$ regardless of the system size, signaling localization of the zero modes near the edges. In contrast, it decreases toward a value $2/L$ in the trivial phase, indicating that the Majorana modes delocalize over the whole lattice on average. The localization of the zero modes in the topological phase is much more visible if we directly look at the spatial profile of the $|[\vec{\psi}_{T}(\bm{s})]_{j}|^{2}$. The inset of Fig. 2(b) shows (trajectory-averaged) values of $|(\vec{\psi}_{T}(\bm{s}))_{j}|^{2}$ for $L=16$ and $\mu_{o}=0.1$. These results indicate that the Lyapunov spectrum clearly reveals edge-localized Majorana zero modes in the topological area-law phase.
Topological invariant and bulk-edge correspondence–
@@
Since $\chi_{T}(\bm{s})$ approximates and converges to $Q_{T}(\bm{s})$ for $T\to\infty$, we will also call $\chi_{T}(\bm{s})$ as the topological invariant. We note that $\widetilde{O}^{\textrm{PBC/APBC}}_{T}(\bm{s})$ is not an orthogonal matrix for general $T$. Note also that the Lyapunov spectrum and topological invariant have been discussed in Refs. [72, 73] for a deterministic nonunitary circuit, while these works did not consider the monitored system determined by the Born rule.
The experimental overhead to compute the topological invariant is exponential in the number of measurements, since the measurement outcomes in the circuit under the APBC are required to be postselected with respect to those in the circuit under the PBC. In general, we have similar postselection problems to observe the measurement-induced phenomena [25]. While there are several works trying to overcome this difficulty [23, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84] and brute-force postselection is possible for small-size systems [85], a definitive method to circumvent postselection problems remains unclear and is reserved for future work.
In Figs. 2(c) and 2(d), we show the lowest single-particle Lyapunov spectrum $z_{1}$ for the measurement-only circuit under the PBC and APBC, respectively. It takes a finite value for both cases, implying a finite bulk Lyapunov gap, within each area-law phase. As shown in Fig. 2(e), the topological invariant $\chi_{T}(\bm{s})$ clearly separates the topological and trivial area-law phases. Note that these results are qualitatively independent of the measurement outcomes $\bm{s}$. In fact, $P^{\mathrm{PBC}}_{T}(\bm{s})=1$ is satisfied in the whole parameter region, since the gap does not close at the transition $\mu_{o}\simeq 0.5$ for the PBC due to finite-size splitting $z_{1}\sim 1/L$. On the other hand, $P^{\mathrm{APBC}}_{T}(\bm{s})$ abruptly changes from $-1$ to $+1$ across the transition point for the APBC, implying an exact gap closing near $\mu_{o}=0.5$. This results in the observed transition in $\chi_{T}(\bm{s})$. The edge-localized zero modes under the OBC in Figs. 2(a) and 2(b) and the topological invariant in Fig. 2(e) indicate that the bulk-edge correspondence applies even to nonequilibrium phases under temporally random measurements, if the bulk gap is open.
Critical gapless phase with additional unitary–
Figure 3: (a) Non-negative single-particle Lyapunov spectrum for the monitored circuit with $J=0.5$ under OBC for $L=128$. (b),(c) Lowest non-negative spectrum $z_{1}$ under (b) PBC and (c) APBC. The inset in (b) shows the system-size dependence of $z_{1}$ at $\mu_{o}=0.1,0.4$, and $0.5$. (d) Topological invariant at $T=L$ averaged over $1000$ different trajectories. The Lyapunov spectrum and topological invariant distinguish the topological area-law, critical sub-volume-law, and trivial area-law phases at $\mu_{o}\simeq 0.4$ and $0.6$, as indicated by the dashed vertical lines.